OpenStudy (apkjv3077):

i need help Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. 1. Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work. i dont understand need someone to walk me through

OpenStudy (kittiwitti1):

What exactly don't you understand? • • • Slope-intercept form is this: \(y=mx+b\) \(m\) is slope \(b\) is the y-intercept

OpenStudy (apkjv3077):

the whole thing i dont even understand the world problem and the question im so confused i even took notes!

OpenStudy (kittiwitti1):

Ok, well, here is the information that is necessary. Some problems tend to overload things to confuse you. http://prntscr.com/btgan1 They want you to convert this equation to slope-intercept form

OpenStudy (apkjv3077):

so the equation is 2x+3y=1,470 i have to put that into slope-intercept form?

OpenStudy (kittiwitti1):

Yes ☺

OpenStudy (darronw):

Hey I remember this from Algebra 1, who's your FLVS teacher?

OpenStudy (apkjv3077):

ok so it would be 3y+2x?

OpenStudy (kittiwitti1):

You have an equation in the form \(Ax+By=C\), now you need to change it to the \(y=mx+b\) form

OpenStudy (kittiwitti1):

Hint: Isolate the \(y\). For example:\[2x+3y=1470\]\[3y=-2x+1470\]This is step one. The \(2x\) is now negative as it was subtracted from both sides, and the left side cancels out.

OpenStudy (apkjv3077):

ok im following that so far so wouldnt i try to get y by its self?

OpenStudy (kittiwitti1):

Yep, can you do it from here? ☺

OpenStudy (apkjv3077):

um maybe lets me try and see where i get too

OpenStudy (kittiwitti1):

Okay, take your time.

OpenStudy (apkjv3077):